OpenStudy (anonymous):

It takes Julia 12 hours to tar a roof, Shayna can tar the same room in 15 hours. Find how long it would take them if they did it together. So far i got 1/x= 1/12+1/15 1/x=15/180+12/180 is this right so far?

OpenStudy (anonymous):

@hero :) :) :) :) :)

hero (hero):

It looks right, but there's a much easier way to solve it.

OpenStudy (anonymous):

Uhh how?

hero (hero):

It is a pretty simple expression called the "working together expression". If two people, J and S, can each complete a task alone, then together they can complete the task in \(\large\frac{J \times S}{J + S}\) amount of time.

OpenStudy (anonymous):

so.. just multiply js on the num and add them on the dem?

hero (hero):

Try it out and see what happens, but make sure you post your result so that I know you are doing it correctly.

OpenStudy (anonymous):

I got 180/27.

hero (hero):

Are you going to divide that to get a more meaningful result?

OpenStudy (anonymous):

6.66?

OpenStudy (anonymous):

like repeating

hero (hero):

6.67 or 6 and 2/3 hours.

hero (hero):

to be exact. Either way, doing it that way is much quicker.

OpenStudy (anonymous):

Okay since its asking to round to the nearest hundredth then i could do 6.67? and also How would i do it if it asked for them to be working alone instead of together?

hero (hero):

Of course, you might want to learn some simplification methods for the fractions.

OpenStudy (anonymous):

think of a clock. what is 2/3 of an hour ?

OpenStudy (anonymous):

Alright and I could do this method for any of the time based problems?

hero (hero):

If it asks, for example, how much did J do alone, then you input everything given into the expression \[\frac{J \times S}{J + S} = t\], then solve for J.

hero (hero):

Suppose you didn't know J, but only knew how much S can do alone, and how much J and S can do together, then you would have: \[\frac{J \times 15}{J + 15} = 6\frac{2}{3}\] Then all you would have to do is simply solve for J by doing some simple mathematical manipulation.

OpenStudy (anonymous):

I'm lost. Okay like my next question is like working together mark and wilbur can clean the attic in 6.3 hours. Had he done it alone it would have taken him 17 hours. How long would it take mark to do it alone. Would i then do 6.3/17?

hero (hero):

By he, they meant Wilbur. and we have to find M

hero (hero):

But anyway to solve: You define your variables as: M = the work Mark can do alone W = the work Wilbur can do alone t = the time they can do the work together You define your working together equation as follows: \[\frac{M \times W}{M + W} = t\] You define your working together equation as follows: \[\frac{M \times 17}{M + 17} = 6.3\]

hero (hero):

So now, all you have to do is just solve for M. Sorry for the confusion.

hero (hero):

By the way, figuring out what to do is just a matter of using the working together equation. All you do is just plug in the appropriate values and solve. But make sure you have the variables properly defined.

OpenStudy (anonymous):

So far i have m17 /m+17= 6.3. So would i add the 17 to the m? or just cross them out

hero (hero):

No, you can't cross anything out. You have to simply figure out how to isolate M. You might want to begin by figuring out how to get rid of the denominator on the left side. By 'get rid of' I don't mean cross anything out. I mean perform the mathematical manipulation necessary to work towards isolating M.

OpenStudy (anonymous):

multiply the 17 by both sides

hero (hero):

No, no. You have to treat the (M + 17) in the denominator as one expression or one number.

OpenStudy (anonymous):

So then it would be two separate problems. mx17= 6.3 and m+17 = 6.3

OpenStudy (anonymous):

hm.

hero (hero):

No it will not. This isn't the zero product property.

OpenStudy (anonymous):

Okay well other than that , I'm completely lost. You can't add the M to the 17. And there both the same so.....it would just cancel out. Other than that guess i don't know the next step

OpenStudy (anonymous):

\[\frac{ M * 17 }{ M + 17 } = 6.3 \rightarrow M * 17 = 6.3(M + 17)\]

OpenStudy (anonymous):

explain...

OpenStudy (anonymous):

just multiply both sides by (M + 17), that will cancel it on the left side

OpenStudy (anonymous):

couldn't you have done that with the addition side too?

hero (hero):

What exactly do you mean by the "addition" side?

OpenStudy (anonymous):

not sure what you mean? in general if \[\frac{ a }{ b } = c\] then \[a = bc\] assuming b is not 0

hero (hero):

There's only two sides to an equation. A right side and a left side. There's no "addition" side.

OpenStudy (anonymous):

okay but how did it go from a fraction to an equation.

OpenStudy (anonymous):

it always was an equation, right? it started as an equation with a fraction on the left hand side, and a single number on the right hand side

hero (hero):

@Artchicky, if you had to solve \[\frac{2x + 4}{5} = 9\] What would you do first?

OpenStudy (anonymous):

I would do the 5 first.

hero (hero):

What would you do with the 5?

OpenStudy (anonymous):

times it by the 2x+4 and the 9

hero (hero):

Okay, and why would you do that? For what purpose?

OpenStudy (anonymous):

to get rid of the fraction

hero (hero):

Okay, so basically, whatever is in the denominator of the fraction, you multiply both sides of the equation by that to get rid of the fraction. So, for the equation we are currently working on, why would we do anything different?

hero (hero):

In order to get rid of the fraction, we multiply both sides of the equation by the whole expression in the denominator, not just part of it, but the whole thing. So not just 17, but M + 17.

OpenStudy (anonymous):

oh because thats whats on the bottom. I get it

hero (hero):

ya, then continue solving from there.

OpenStudy (anonymous):

17m= 6.3m+ 107.1?

hero (hero):

Okay, continue solving for m. You have to place like terms on the same side, right? I'm pretty sure you know how to do that.

OpenStudy (anonymous):

10.0

hero (hero):

You mean M = 10. That is correct.

hero (hero):

This wasn't as difficult as you made it out to be.

OpenStudy (anonymous):

well it went from an equation to a problem with two people putting out two different problems.

hero (hero):

There's more than one way to solve math problems. That's the general consensus. I was trying to show you a method that was supposed to be an easier, simpler method. However, for some reason, you seemed to have trouble solving the fractional equation, with the most confusing thing for you being how to properly isolate the variable M. But hopefully you have it now.

OpenStudy (anonymous):

Yes i understand it

hero (hero):

The method I was showing you was a three step process: 1. Get rid of the denominator 2. Distribute 3. Isolate the variable The method saifoo.khan showed you involves finding LCDs and such, which is fine, but it is always good to have an alternative method.

hero (hero):

The working together expression works best when you know how much time both of a job persons can finish alone and need to find how much time it takes if they work together.